OpenStudy (anonymous):

Sum of series help please: How do you calculate the sum of: .4+.6^2(.4)+.6^4(.4)+.6^6(.4)...

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{ this is the series ?} \hspace{.33em}\\~\\ 0.4+0.6^2(0.4)+0.6^4(0.4)+0.6^6(0.4)... \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

hint: \(6^2=.36\)

OpenStudy (misty1212):

sum an normal \[\sum ar^n=\frac{a}{1-r}\] with \[a=.4,r=.36\]

OpenStudy (anonymous):

Hi @mathmath333, yes, that is the series.

OpenStudy (mathmath333):

did u understand misty's solution